OpenStudy (anonymous):

6^(x+1)=216 I have absolutely no idea how to solve this..

ganeshie8 (ganeshie8):

Hint : \[\large 216 = 6^3\]

ganeshie8 (ganeshie8):

\[\large 6^{x+1} = 6^3\] since the bases are equal, exponents also must be equal : \[\large x+1 = 3\] solve \(\large x\)